OpenStudy (liv33):

Medal and fan to simplify: -√(121/16)

OpenStudy (kinged):

\[\sqrt{\frac{ x }{ y }}=\frac{ \sqrt{x} }{ \sqrt{y} }\]

OpenStudy (liv33):

Could you elaborate?

OpenStudy (kinged):

\[-\sqrt{\frac{ 121 }{ 16 }}=-\frac{ \sqrt{121} }{ \sqrt{16} }\]